OpenStudy (howard-wolowitz):

Am I right : I got 0.95339702

OpenStudy (anonymous):

what was the last step you did to solve this problem

OpenStudy (howard-wolowitz):

@wolf1728 I got 0.9533

OpenStudy (wolf1728):

9^2x = 66 2x * log(9) = log(66) 2x = log(66) / log(9) 2x = 1.9067940461 x = 0.9533970231 Seems I get .9533 also

OpenStudy (mrnood):

YUP - me too